# Efficient Low-Memory Fast Stack Decoding with Variance Polarization for PAC Codes

### v1

@@ Abstract
Polarization-adjusted convolutional (PAC) codes have recently emerged as a promising class of error-correcting codes, achieving near-capacity performance particularly in the short block-length regime. In this paper, we propose an enhanced stack decoding algorithm for PAC codes that significantly improves parallelization by exploiting specialized bit nodes, such as rate-0 and rate-1 nodes. For a rate-1 node with $N_{0}$ leaf nodes in its corresponding subtree, conventional stack decoding must either explore all $2^{N_{0}}$ paths, or, same as in fast list decoding, restrict attention to a constant number of candidate paths. In contrast, our approach introduces a pruning technique that discards wrong paths with a probability exponentially approaching zero, retaining only those whose path metrics remain close to their expected mean values. Furthermore, we propose a novel approximation method for estimating variance polarization under the binary-input additive white Gaussian noise (BI-AWGN) channel. Leveraging these approximations, we develop an efficient stack-pruning strategy that selectively preserves decoding paths whose bit-metric values align with their expected means. This targeted pruning substantially reduces the number of active paths in the stack, thereby decreasing both decoding latency and computational complexity. Numerical results demonstrate that for a PAC(128,64)(128,64) code, our method achieves up to a 70% reduction in the average number of paths without degrading error-correction performance.
Index Terms:
PAC codes, polar codes, stack decoding, varentropy, polarization, mutual information, sequential decoding.
@@ I Introduction
As a sequential decoding algorithm, the stack algorithm [5, 6] evaluates the most likely path in the stack, computes the metric values for both of its branches, and adds them to the stack after removing that path. Since the stack algorithm can revisit certain levels of the decoding tree multiple times, adding both branches at each step can greatly increase the number of paths in the stack, complicating the identification of the most likely path and increasing the decoding delay. In this paper, we introduce a stack-pruning method that identifies bifurcated branches unlikely to belong to the correct path, thus preventing their addition to the stack. Our results demonstrate that a pruning approach significantly reduces the average number of paths in the stack, thereby improving both decoding complexity and delay.
Our stack-pruning technique benefits from the concept of metric function polarization introduced in [7], where it is shown that at each step of the polarization process, the mean and variance of the metric function random variable correspond to the polarized mutual information and varentropy of the synthesized binary input discrete memoryless channel (BI-DMC), respectively. For a binary input additive white Gaussian noise (BI-AWGN) channel, [8] proposes an algorithm to calculate the polarized mutual information, which can also be used to compute the polarized channel varentropy [7]. In [7], also an approximation method for calculating the varentropy of the BI-AWGN channel is introduced. Following the approach in [9], in this paper, we also assume that the channel remains a BI-AWGN channel after each polarization step. Using the method outlined in [7], we propose a technique to compute the polarized varentropies of the synthesized BI-AWGN channels at each step of the polarization. Numerical results show that the polarized varentropy values obtained from our approximation method closely match those derived using the method from [8].
@@
To enhance the parallelism of our proposed stack-pruning algorithm, we leverage four types of special nodes: rate-0, rate-1, rate-IV and repetition (REP) nodes. A rate-1 node with $N_{0}$ leaf nodes can generate $2^{N_{0}}$ candidate paths in the conventional stack decoding algorithm. In [11], a fast block stack decoding method for polar codes is introduced, which recommends retaining only 4 out of the $2^{N_{0}}$ possible paths to avoid significant ECP loss. In polar codes, the rate profile is typically constructed based on the reliability of the polarized bit-channels, and retaining 4 paths generally preserves ECP performance. However, for short block-length PAC codes, a rate profile based on alternative criteria—such as the weight distribution of the code—is often more favorable [3]. In our approach, for rate-1 nodes, we apply a stack-pruning strategy that permits bifurcation only when the corresponding child node has a bit metric exceeding our proposed threshold. For REP nodes, which yield two candidate paths, we retain a path only if the bit metric of its final leaf node exceeds its associated threshold. Our proposed algorithm for handling rate-IV and repetition (REP) special nodes can also be seamlessly applied to any other rate profile regardless of the particular strategy to select information-bit positions. Such nodes are investigated in the context of SCL decoding in [12]. In particular, since the pruning and metric-update rules we introduce do not depend on a particular rate profile structure, the method remains applicable even when the bit-channel ordering or information set is tailored to alternative design criteria. Our contributions in this paper can be summarized as follows:
@@
The remainder of this paper is organized as follows. Section II reviews the necessary background and preliminaries to make the paper self-contained. In Section III, we discuss the polarization of channel parameters and introduce an approximation method for estimating the polarized varentropy. Section IV presents our proposed pruning strategy, which selectively removes unlikely decoding paths from the stack based on adaptive thresholds. In Section V, we introduce the fast stack decoding algorithm, which efficiently handles different types of special nodes and integrates the proposed pruning method to reduce decoding complexity. Finally, Section VI provides numerical results demonstrating the performance gains and computational benefits of the proposed techniques.
@@ II-B PAC Coding Scheme
II-C Channel Parameters
@@ II-E Path Metric
II-F BI-AWGN
For a BI-AWGN channel, the channel cutoff rate is given by [14, p. 142]
@@ III Polarization of Channel Parameters
Following the approach in [9], we assume that the log-likelihood ratio (LLR) of the channel output $W$ follows a Gaussian distribution with mean $2/\sigma^{2}$ and variance $4/\sigma^{2}$, and that after each polarization step, the resulting channels continue to behave as BI-AWGN channels. This allows the computation of $\sigma_{N}^{(i)}$ after $\log_{2}(N)$ polarization steps for the $i$th polarized channel. Building on this model, in this paper we use the cutoff rate of the $i$th bit-channel as
@@
Similarly, the $i$th bit-channel mutual information $I(W_{N}^{(i)})$ can be approximated by $J_{\text{approx}}(t)$, and the bit-channel varentropy can be approximated by $V_{\text{approx}}(t)$, where $t=\frac{2}{\sigma_{N}^{(i)}}$.
@@ IV Stack pruning method
[7]. In a BI-DMC, the bit-metric random variable $\phi(U_{i};\mathbf{Y},\mathbf{U}^{i-1})$ converges to its expected value in probability. More precisely, it holds that
@@
Moreover, using Chernoff’s bound, the following theorem from [10] shows that the probability of the bit-metric random variable deviating significantly from its expected value decays exponentially to zero.
@@
Utilizing the upper bounds from (14) and (15), we propose a path-pruning strategy for the stack decoding algorithm as follows. If the bit-metric value $\gamma(u_{i};\mathbf{y},\mathbf{u}^{i-1})$ corresponding to a data bit is less than $\lfloor-(\mathrm{Var}(W_{N}^{(i)})/P_{th})^{1/2}\,\rfloor-10$ or less than $\lfloor\log_{2}(P_{th})\rfloor$, then the corresponding branch is not explored and is not inserted into the stack. We define a pruning threshold vector $\boldsymbol{\gamma}_{T}$, whose $i$th element is given by
$\boldsymbol{\gamma}_{T}^{(i)}=\min\left\{\left\lfloor-\sqrt{\frac{\mathrm{Var}(W_{N}^{(i)})}{P_{th}}}\right\rfloor-10,\;\left\lfloor\log_{2}(P_{th})\right\rfloor\right\}.$
@@ V Fast stack algorithm
We employ a heap data structure [16, 17], which enables logarithmic-time operations, to implement a memory-efficient stack decoding algorithm for PAC codes. In this structure, partial path metrics serve as the keys, while the satellite data consists of four components: the convolutional code’s input vector, output vector, and state; and the polar demapper’s intermediate log-likelihood ratios (LLRs). The intermediate LLRs and the convolutional code’s input/output vectors are stored in arrays of length N1N-1 and $N$, respectively. For an alternative implementation that stores checksum values (instead of the convolutional code’s outputs) in a vector of length $N/2$, see [18].
@@
28:    $S,S_{N}]\leftarrow\textsc{f-REP}(\textit{topS},S,S_{N},\mathbf{st},\mathbf{L},\mathbf{idx}_{\text{LLR}},\mathbf{idx}_{\text{chunk}},$
29:        $\mathbf{u},\mathbf{c},\mathbf{E}_{0},\mathbf{RP},\boldsymbol{\gamma}_{T},\text{S-Size},\mathbf{r}_{i},N_{\text{chunk}})$
30:  else if $\textit{Type}=1$ then
31:    $S,S_{N}]\leftarrow\textsc{f-Type-1}(\textit{topS},S,S_{N},\mathbf{st},\mathbf{L},\mathbf{idx}_{\text{LLR}},$
32:        $\mathbf{idx}_{\text{chunk}},\mathbf{u},\mathbf{c},\mathbf{E}_{0},\boldsymbol{\gamma}_{T},\text{S-Size},\mathbf{r}_{i},N_{\text{chunk}})$
33:  else if $\textit{Type}=\text{IV}$ then
34:    $S,S_{N}]\leftarrow\textsc{f-Type-IV}(\textit{topS},S,S_{N},\mathbf{st},\mathbf{L},\mathbf{idx}_{\text{LLR}},$
35:        $\mathbf{idx}_{\text{chunk}},\mathbf{u},\mathbf{c},\mathbf{E}_{0},\mathbf{RP},\boldsymbol{\gamma}_{T},\text{S-Size},\mathbf{r}_{i},N_{\text{chunk}})$
36:  else if $\textit{Type}=\text{ALL}$ then
37:    $S,S_{N}]\leftarrow\textsc{f-Type-All}(\textit{topS},S,S_{N},\mathbf{st},\mathbf{L},\mathbf{idx}_{\text{LLR}},$
38:        $\mathbf{idx}_{\text{chunk}},\mathbf{u},\mathbf{c},\mathbf{E}_{0},\boldsymbol{\gamma}_{T},\text{S-Size},\mathbf{r}_{i},N_{\text{chunk}})$
@@
40:  if $\text{length}(S[1].\mathbf{v})=N$ then
41:    break
42:  end if
43:end while
@@
46:return $\hat{\mathbf{v}},\textit{cycles},S\textit{-used}]$
Algorithm 2 presents the main function of our proposed fast stack decoding algorithm for PAC codes. The algorithm operates on the channel output LLRs, denoted by $\mathbf{L}_{ch}$, the data index set $\mathcal{A}$, the connection polynomial $\mathbf{c}$, the vector of bit-channel bias values (bit-channel cutoff rates) $\mathbf{E}_{0}$, the maximum number of decoding cycles, the stack size, and the threshold vector $\boldsymbol{\gamma}_{T}$, which determines whether a path should be pruned. The outputs of the algorithm are the data carrier vector $\hat{\mathbf{v}}$, the number of cycles, and the stack size used (i.e., the number of paths stored in the stack).
@@
6:  $\mathbf{st}\leftarrow[0,~\mathbf{st}[1:\,c-1]]$
@@
Algorithm 4 f-REP (Rate $1/N_{\text{chunk}}$ Node)
@@
13:  $\mathbf{ST}\leftarrow\big{[}\mathbf{V}_{0}[:][t],~\mathbf{ST}[:][1:c-1]\big{]}$
@@
16:  $\mathbf{ST}\leftarrow\big{[}\mathbf{V}_{0}[:][t],~\mathbf{ST}[:][1:c-1]\big{]}$
@@
51:    $\mathbf{st}^{\text{temp}}\leftarrow[\,\mathbf{V}_{0}[t],~\mathbf{st}^{\text{temp}}[1:c-1]\,]$

### v2

@@ Abstract
Polarization-adjusted convolutional (PAC) codes have recently emerged as a promising class of error-correcting codes, achieving near-capacity performance particularly in the short block-length regime. In this paper, we propose an enhanced stack decoding algorithm for PAC codes that significantly improves parallelization by exploiting specialized bit nodes, such as rate-0 and rate-1 nodes. For a rate-1 node with $N_{0}$ leaf nodes in its corresponding subtree, conventional stack decoding must either explore all $2^{N_{0}}$ paths, or, same as in fast list decoding, restrict attention to a constant number of candidate paths. In contrast, our approach introduces a pruning technique that removes candidate paths with small path metrics while ensuring that the probability of pruning the correct path decays exponentially with the threshold. Furthermore, we propose a novel approximation method for estimating variance polarization under the binary-input additive white Gaussian noise (BI-AWGN) channel. Leveraging these approximations, we develop an efficient stack-pruning strategy that selectively preserves decoding paths whose bit-metric values align with their expected means. This targeted pruning substantially reduces the number of active paths in the stack, thereby decreasing both decoding latency and computational complexity. Numerical results demonstrate that for a PAC(128,64)(128,64) code, our method achieves up to a 70% reduction in the average number of paths without degrading error-correction performance.
@@ I Introduction
As a sequential decoding algorithm, the stack algorithm was first introduced for convolutional codes [35, 8]. At each step, it selects the most likely path in the stack, extends it along both branches, computes the corresponding path metrics, and inserts the resulting paths into the stack after removing the original path. Stack decoding and related sequential tree-search methods have also been studied for both polar and PAC codes [22, 26, 31, 7, 28, 19, 24]. Because the stack algorithm may revisit the same decoding levels multiple times, extending both branches at each step can significantly increase the number of candidate paths in the stack, which complicates the search for the most likely path and increases decoding delay. In this paper, we propose a stack-pruning method that identifies branches unlikely to lie on the correct path and prevents them from being inserted into the stack. Our results show that the proposed pruning approach significantly reduces the average stack size, thereby reducing both decoding complexity and delay.
Our stack-pruning technique benefits from the concept of metric function polarization introduced in [13], where it is shown that at each step of the polarization process, the mean and variance of the metric function random variable correspond to the polarized mutual information and varentropy (variance) of the synthesized binary input discrete memoryless channel (BI-DMC), respectively. For a binary input additive white Gaussian noise (BI-AWGN) channel, [27] proposes an algorithm to calculate the polarized mutual information, which can also be used to compute the polarized channel varentropy [13]. In [13], also an approximation method for calculating the varentropy of the BI-AWGN channel is introduced. Following the approach in [11], in this paper, we also assume that the channel remains a BI-AWGN channel after each polarization step. Using the method outlined in [13], we propose a technique to compute the polarized varentropies of the synthesized BI-AWGN channels at each step of the polarization. Numerical results show that the polarized varentropy values obtained from our approximation method closely match those derived using the method from [27].
@@
To enhance the parallelism of the proposed fast stack decoder, we consider four types of special nodes: rate-0, rate-1, type-IV, and REP nodes. Other special nodes, such as single-parity-check (SPC) [34, 33, 9] and Type-T nodes [6], are also of interest, and extending the proposed framework to incorporate them is left for future work. A rate-1 node with $N_{0}$ leaf nodes can generate $2^{N_{0}}$ candidate paths in the conventional stack decoding algorithm. In [29], a fast block stack decoding method for polar codes is introduced, which recommends retaining only 4 out of the $2^{N_{0}}$ possible paths to avoid significant ECP loss. In polar codes, the rate profile is typically constructed based on the reliability of the polarized bit-channels, and retaining 4 paths is generally sufficient to preserve ECP. However, for short block-length PAC codes, a rate profile based on alternative criteria—such as the weight distribution of the code—is often more favorable [15]. In our approach, for rate-1 nodes, we apply a stack-pruning strategy that permits bifurcation only when the corresponding child node has a bit metric exceeding our proposed threshold. For REP nodes, which yield two candidate paths, we retain a path only if the bit metric of its final leaf node exceeds its associated threshold. Our proposed algorithm for handling rate-IV and REP special nodes can also be seamlessly applied to any other rate profile regardless of the particular strategy to select information-bit positions. Such nodes are investigated in the context of successive-cancellation list (SCL) decoding in [32]. In particular, since the pruning and metric-update rules we introduce do not depend on a particular rate profile structure, the method remains applicable even when the bit-channel ordering or information set is tailored to alternative design criteria. Our contributions in this paper can be summarized as follows:
@@
The remainder of this paper is organized as follows. Section II reviews the necessary background and preliminaries to make the paper self-contained. In Section III, we discuss the polarization of channel parameters and introduce an approximation method for estimating the polarized varentropy. Section IV presents our proposed pruning strategy, which selectively removes unlikely decoding paths from the stack based on adaptive thresholds. In Section V, we introduce the fast stack decoding algorithm, which efficiently handles different types of special nodes and integrates the proposed pruning method to reduce decoding complexity. Section VI provides numerical results demonstrating the performance gains and computational benefits of the proposed techniques. Finally, Section VII concludes the paper.
@@ II-B PAC Coding Scheme
II-C Rate Profile
For short-blocklength PAC codes, the choice of the rate profile $\mathcal{A}$ has a significant impact on both the ECP and the decoding complexity. As shown in [3, 15], PAC codes with a Reed–Muller (RM)-based rate profile can achieve performance close to theoretical bounds for certain code lengths and code rates. In an RM-based construction, the information set $\mathcal{A}$ is selected according to the rows of $\mathbf{F}^{\otimes n}$ with the largest Hamming weights. A limitation of a pure RM-based construction is that it is not possible for designing PAC codes at arbitrary rates. To obtain rate profiles for a broader range of rates, alternative construction methods can be employed. One approach is to use a genetic-algorithm-based optimization initialized from the RM profile, which enables the design of rate profiles for additional rates [14]. Another flexible approach is Monte-Carlo (MC)-based construction, which can be used to design rate profiles for essentially arbitrary rates [17]. Reinforcement-learning-based methods (Q-learning) have also been proposed for rate-profile optimization [12]. In this work, the proposed fast stack decoding and pruning framework is not restricted to a specific rate-profile construction. Once the information set $\mathcal{A}$ is specified, the same decoding algorithm can be applied to RM-based, Q-learning, genetic-algorithm-based, and MC-based rate profiles.
II-D Channel Parameters
@@ II-F Path Metric
II-G BI-AWGN
For a BI-AWGN channel, the channel cutoff rate is given by [30, p. 142]
@@ II-G BI-AWGN
II-H Sequential Decoders for PAC Codes
Stack decoding and Fano decoding are both sequential decoders, but they offer different and practically important tradeoffs. In stack decoding, the decoder always expands the currently most promising partial path (highest metric) and keeps competing candidates available for later refinement. This best-first behavior avoids long excursions on weak paths and reduces the sensitivity to step-size/threshold-tuning choices that are inherent in Fano decoding search. In particular, Fano decoding relies on a threshold rule: when forward progress along the current path fails to meet the threshold, the decoder decreases the threshold and/or backtracks to earlier depths to explore alternative branches. Such backtracking can be frequent in low-SNR regimes or for unfavorable local metric fluctuations, creating highly variable decoding time and additional latency due to repeated revisits of previously explored tree levels. Moreover, stack decoding provides an explicit complexity/latency knob: by limiting the stack size and the number of cycles, one can enforce a predictable worst-case budget while retaining near-optimal performance. In our experiments, a bounded stack decoder achieves FER comparable to Fano decoding even when Fano is allowed to explore an unbounded number of paths (see Fig. 2 and Fig. 3).
Compared to list decoding (e.g., SCL), stack decoding can be viewed as an adaptive list method: instead of maintaining a fixed list of size $L$ at every information bit, the stack keeps only those partial paths that remain competitive under the path-metric ordering. This is particularly attractive for short-blocklength PAC codes, where high-rate segments (e.g., rate-1 subtrees) can otherwise trigger a large expansion; classical fast-list approaches must restrict the candidate set to a small constant to control complexity, which may not be robust across PAC rate profiles. In contrast, our proposed pruning strategy leverages metric polarization/varentropy to admit bifurcations only when the corresponding bit-metrics are consistent with their expected behavior, reducing the average number of active paths substantially (up to $\approx 70\%$ in our results) without degrading error-correction performance. Finally, stack decoding admits an efficient implementation via a heap-based data structure with logarithmic-time stack operations, enabling a low-memory decoder that naturally carries the per-path convolutional state needed for PAC decoding.
@@ III Polarization of Channel Parameters
Following the approach in [11], we assume that the log-likelihood ratio (LLR) of the channel output $W$ follows a Gaussian distribution with mean $2/\sigma^{2}$ and variance $4/\sigma^{2}$, and that after each polarization step, the resulting channels continue to behave as BI-AWGN channels. This allows the computation of $\sigma_{N}^{(i)}$ after $\log_{2}(N)$ polarization steps for the $i$th polarized channel. Building on this model, in this paper we propose using the cutoff rate of the $i$th bit-channel as
@@
Similarly, we propose that the mutual information of the polarized $i$th bit-channel, $I(W_{N}^{(i)})$, be approximated by $J_{\text{approx}}(t)$, and that its varentropy be approximated by $V_{\text{approx}}(t)$, where $t=\frac{2}{\sigma_{N}^{(i)}}$.
@@ IV Stack pruning method
[13]. In a BI-DMC, the bit-metric random variable $\gamma(U_{i};\mathbf{Y},\mathbf{U}^{i-1})$ converges to its expected value in probability. More precisely, it holds that
@@
Note that, in proving Theorem 2, the blocklength must tend to infinity. In that asymptotic regime, the upper bound in (14) converges to zero a.e. However, PAC codes are mainly studied in the short-blocklength regime, where the variances of the polarized bit-channels are not yet fully polarized. In contrast, the bound in Theorem 1 is non-asymptotic and is therefore useful for short blocklengths, providing more insight into the finite-length behavior than Theorem 2.
Moreover, using Chernoff’s bound, the following theorem from [16] shows that the probability of the bit-metric random variable deviating significantly from its expected value decays exponentially to zero. Similarly, this theorem is also useful for short-blocklength PAC codes.
@@
Utilizing the upper bounds from (14) and (15), we propose a path-pruning strategy for the stack decoding algorithm as follows. Let $P_{\mathrm{th}}$ denote a design parameter specifying the target upper bound on the probability of pruning the correct branch. By inverting the Chebyshev and Chernoff bounds, we obtain two candidate thresholds for the $i$th bit-channel:
$\gamma_{T,\mathrm{Cheb}}^{(i)}=\left\lfloor-\sqrt{\frac{\mathrm{Var}(W_{N}^{(i)})}{P_{\mathrm{th}}}}\right\rfloor-10,~~\gamma_{T,\mathrm{Cher}}^{(i)}=\left\lfloor\log_{2}(P_{\mathrm{th}})\right\rfloor.$
We then define a pruning threshold vector $\boldsymbol{\gamma}_{T}$, whose $i$th element is
$\boldsymbol{\gamma}_{T}^{(i)}=\min\left\{\gamma_{T,\mathrm{Cheb}}^{(i)},\gamma_{T,\mathrm{Cher}}^{(i)}\right\}.$
Accordingly, if the bit-metric value $\gamma(u_{i};\mathbf{y},\mathbf{u}^{i-1})$ corresponding to a data bit is less than $\boldsymbol{\gamma}_{T}^{(i)}$, then the corresponding branch is not explored and is not inserted into the stack. We use an additional constant offset in the Chebyshev-based threshold as an empirical safety margin.
@@ V Fast stack algorithm
We employ a heap data structure [5, 25], which enables logarithmic-time operations, to implement a memory-efficient stack decoding algorithm for PAC codes. In this structure, partial path metrics serve as the keys, while the satellite data consists of four components: the convolutional code’s input vector, output vector, and state; and the polar demapper’s intermediate LLRs. The intermediate LLRs and the convolutional code’s input/output vectors are stored in arrays of length N1N-1 and $N$, respectively. For an alternative implementation that stores checksum values (instead of the convolutional code’s outputs) in a vector of length $N/2$, see [20].
@@
28:    $S,S_{N}]\leftarrow\textsc{f-Type-REP}(\textit{topS},S,S_{N},\mathbf{st},\mathbf{L},\mathbf{idx}_{\text{LLR}},$
29:        $\mathbf{idx}_{\text{chunk}},\mathbf{u},\mathbf{c},\mathbf{E}_{0},\mathbf{RP},\boldsymbol{\gamma}_{T},\text{S-Size},\mathbf{r}_{i},N_{\text{chunk}})$
30:  else if $\textit{Type}=\text{IV}$ then
31:    $S,S_{N}]\leftarrow\textsc{f-Type-IV}(\textit{topS},S,S_{N},\mathbf{st},\mathbf{L},\mathbf{idx}_{\text{LLR}},$
32:        $\mathbf{idx}_{\text{chunk}},\mathbf{u},\mathbf{c},\mathbf{E}_{0},\mathbf{RP},\boldsymbol{\gamma}_{T},\text{S-Size},\mathbf{r}_{i},N_{\text{chunk}})$
33:  else if $\textit{Type}=\text{ALL}$ then
34:    $S,S_{N}]\leftarrow\textsc{f-Type-All}(\textit{topS},S,S_{N},\mathbf{st},\mathbf{L},\mathbf{idx}_{\text{LLR}},$
35:        $\mathbf{idx}_{\text{chunk}},\mathbf{u},\mathbf{c},\mathbf{E}_{0},\boldsymbol{\gamma}_{T},\text{S-Size},\mathbf{r}_{i},N_{\text{chunk}})$
36:  end if
37:  if $\text{length}(S[1].\mathbf{v})=N$ then
38:    break
@@
40:end while
@@
43:return $\hat{\mathbf{v}},\textit{cycles},S\textit{-used}]$
Algorithm 2 presents the main function of our proposed fast stack decoding algorithm for PAC codes. The algorithm operates on the channel output LLRs, denoted by $\mathbf{L}_{ch}$, the data index set $\mathcal{A}$, the binary coefficient vector $\mathbf{c}$ corresponding to the connection polynomial $c(t)$, where the convolutional memory depth is $|\mathbf{c}|-1$, the vector of bit-channel bias values (bit-channel cutoff rates) $\mathbf{E}_{0}$, the maximum number of decoding cycles, the stack size, and the threshold vector $\boldsymbol{\gamma}_{T}$, which determines whether a path should be pruned. The outputs of the algorithm are the data carrier vector $\hat{\mathbf{v}}$, the number of cycles, and the stack size used (i.e., the number of paths stored in the stack).
@@
6:  $\mathbf{st}\leftarrow[0,~\mathbf{st}[1:\,|\mathbf{c}|-1]]$
@@
Algorithm 4 f-Type-REP (Rate $1/N_{\text{chunk}}$ Node)
@@
13:  $\mathbf{ST}\leftarrow\big[\mathbf{V}_{0}[:][t],~\mathbf{ST}[:][1:|\mathbf{c}|-1]\big]$
@@
16:  $\mathbf{ST}\leftarrow\big[\mathbf{V}_{0}[:][t],~\mathbf{ST}[:][1:|\mathbf{c}|-1]\big]$
@@
51:   $\mathbf{st}^{\text{temp}}\leftarrow[\,\mathbf{V}_{0}[t],~\mathbf{st}^{\text{temp}}[1:|\mathbf{c}|-1]\,]$